OpenStudy (anonymous):

Please help!!!!! MEDAL AND FAN!!!!! Referring to the Fig. in Question #8, segment AB and segment NM are diameters of circle E. Find the angle of arc MCB.

OpenStudy (anonymous):

This is the picture

Nnesha (nnesha):

 that straight line add up to 1180 degree right so if what would you add into 55+65 to get 180 ?

Nnesha (nnesha):

180**

OpenStudy (anonymous):

I would have to add 60 I think

Nnesha (nnesha):

you think? 65+60+55= 180 ?

OpenStudy (anonymous):

Yep..... because 55+65=120+60=180

Nnesha (nnesha):

 that middle angle is equal to arc length and remember circle add up to 360 degree